# A template for the arxiv style

### v1

@@ Abstract
We consider the task of identifying the causal parents of a target variable among a set of candidate variables from observational data. Our main assumption is that the candidate variables are observed in different environments which may, for example, correspond to different settings of a machine or different time intervals in a dynamical process. Under certain assumptions different environments can be regarded as interventions on the observed system. We assume a linear relationship between target and covariates, which can be different in each environment with the only restriction that the causal structure is invariant across environments. This is an extension of the ICP (Invariant Causal Prediction) principle by Peters et al. (2016), who assumed a fixed linear relationship across all environments. Within our proposed setting we provide sufficient conditions for identifiability of the causal parents and introduce a practical method called LoLICaP (Locally Linear Invariant Causal Prediction), which is based on a hypothesis test for parent identification using a ratio of minimum and maximum statistics. We then show in a simplified setting that the statistical power of LoLICaP converges exponentially fast in the sample size, and finally we analyze the behavior of LoLICaP experimentally in more general settings.
@@ 1 Introduction
We consider the problem of identifying the causal parents of a target variable, among a set of candidate variables, based only on observational data. The main assumption we make is that data is observed in different environmental scenarios. An emblematic example is that of machine diagnostics, where we are monitoring several connected components of the machine. Different environments correspond, for example, to machines in different locations, different settings of a machine, or data collected in different points in time. If the system behaves differently across environments we talk about heterogeneous environments, and these can be interpreted as accidental interventions on the system. This means that, under certain modeling assumptions, one can leverage the information provided by those accidental interventions to perform causal discovery tasks. One particular way of using this information is based on the idea that the performance of any prediction model for the target variable should be invariant under interventions (in different environments) if and only if this prediction model uses exclusively the causal parents of the target: For example, a model that uses the color green to predict that an image contains a cow may work well if we only have cow pictures from grassy areas. The green background is, however, only a correlated variable and not a causal parent, and the performance of such a model should significantly drop if we suddenly present it with cow images with a sandy background. This principle of invariant causal prediction (ICP) was proposed by Peters et al. (2016), who assumed that the causal relationship between the target and its parents follows a linear model which is fixed across all environments. We extend upon that work by relaxing the global linearity assumption to a locally linearity assumption, meaning that each environment is equipped with its own linear model. Relaxing the global model to local models carries a couple of interesting implications that we address in this paper. In particular, locally linear models extend the scope of what might be considered heterogeneous environments. Heterogeneity as viewed in Peters et al. (2016) is always introduced by changes in the feature distributions. Considering locally linear models, heterogeneity can also be induced by parameter changes in the local models, which can be seen as changes in the causal mechanism between different environments. The main contributions of this paper are as follows:
  • We extend the work of Peters et al. (2016) to a locally linear model class.
  • We provide sufficient conditions under which the causal parents of a target variable are identifiable within the model class under investigation.
  • Given a finite amount of data, we introduce LoLICaP (Locally Linear Invariant Causal Prediction), a method suitable to decide if a given set of potential causal parents is a plausible causal explanation of a target variable.
  • Under certain normality assumptions we show that the power of LoLICaP converges exponentially fast, in the sample size, towards one.
  • We provide experimental evidence that LoLICaP is applicable in a broad scope of settings.
Organization:
In Section 2 we contextualize our contributions by discussing related work. In Section 3 we introduce the setting, our assumptions and formalize inference goals. This section features, in addition, concrete examples illustrating the meaningfulness of the modeling assumptions. In Section 4 we present a meta-approach and give a characterization of its potential in terms of identifying causal relations. Motivated by that, Section 5 presents a concrete approach to detect causal parents with limited amount of data. Section 6 illustrates the performance of LoLICaP in a variety of settings. We conclude the paper in Section 7 with a discussion and an outlook on future work.
@@ 2 Related work
The idea of relying on heterogeneous environments for causal discovery is not new, and was instrumental for the work in Peters et al. (2016). This setup was also extended to sequential data in Pfister et al. (2019), while Heinze-Deml et al. (2018) investigate non-linear models by using conditional independence tests. Some papers relax the global linearity assumption for local models with the help of structural assumptions between the local models: Christiansen and Peters (2020) and Zhou et al. (2022) assume that the local structural parameters, i.e. the parameters of the assumed linear relationship between target and its parents, are related through a common (unobserved) variable. Huang et al. (2019) considers a temporal setting, where the local structural parameters change according to an auto-regressive model. The same authors extent this to a constraint based method under the assumption that the changes in the parameters happen independently (Huang et al., 2020). In contrast to these works we consider a setting where the structural parameters can be radically different for the various environments.
The concept of causal invariance is also carried to the context of out-of-domain generalization for machine learning models (Arjovsky et al., 2019; Oberst et al., 2021; Wald et al., 2021). The main underlying idea is that causal features, in contrast to merely correlated features, should be useful in all domains. In other words, the predictive power of causal features should be invariant in all domains. Differences in the approaches in that literature often stem from which parts of the model are assumed to be invariant across environments. While the goal of out-of-domain generalization is different than our goal, which is causal discovery, many of the concepts and models resemble each other.
The following papers aggregate the information of different environments, without needing the heterogeneity aspect, while typically assuming that the causal structure is acyclic, and thus representable by a directed acyclic graph (DAG). Osman et al. (2023) propose a method that finds causal structures and detects interventions using a minimum description length score on the causal factorization, which they define jointly over the different environments. They assume non-linear relationships, identifiability of the DAG up the Markov equivalence class and a low noise condition. Chen et al. (2021) assume a linear structural equation model (SEM), such that at least in some environments the corresponding DAG is identifiable from data. To find a solution of the DAG they propose a penalized maximum likelihood estimator which is optimized jointly over all environments. Shimizu (2012) proposes an extension of LiNGAM (Shimizu et al., 2006) to the multiple environment case. LiNGAM is a method that can find DAGs under a linearity and non-Gaussianity assumption.
Finally, Mooij et al. (2020) developed the joint causal inference (JCI) framework, where environments (in the paper called contexts) are directly encoded as part of the structural causal model (SCM). Many existing methods, in particular also the work of Peters et al. (2016), can be stated in that framework, and many methods for single environments, e.g. fast causal inference by Spirtes (2001), can be used within the JCI framework.
@@ 3 Setting
In this work we consider a scenario where there are two observable quantities of interest: a target, denoted by $Y$, and a set of covariates $X\vcentcolon=(X_{1},\cdots,X_{D})$ for $D\in\mathbb{N}$. Our overarching goal is to identify which of the $D$ covariates are the causal parents of the target $Y$. What we precisely mean with a causal parent is detailed in the next subsection. To specify the relation we assume between target and covariates, we introduce the notion of an environment. Specifically, suppose we have $E\in\mathbb{N}$ different environments and in each environment we receive $n^{e}$ observations, so that $Y^{e}\in\mathbb{R}^{n^{e}}$ and $X^{e}\in\mathbb{R}^{n^{e}\times D}$ are respectively all the target and covariate observations. Our main assumption is that for each $e\in[E]$ there exists a column vector $\beta^{e}\in\mathbb{R}^{D}$ such that
$Y^{e}=X^{e}\beta^{e}+\varepsilon^{e},$ (1)
where $\varepsilon^{e}\in\mathbb{R}^{n^{e}}$ is a zero-mean random noise. The specific independence assumptions on $\varepsilon^{e}$ are made explicit below. In this work we consider the assumed relationship $Y^{e}=X^{e}\beta^{e}+\varepsilon^{e}$ to be a structural causal model (SCM) in the sense of Pearl (2016), and thus we consider the set $S^{e,*}\vcentcolon=\{d\in[D]\mid\beta^{e}_{d}\neq 0\}\subseteq[D]$ as the true causal parents of $Y^{e}$. This implies that $Y^{e}$ can have different causal parents in each environment, and we consider a variable with index $d\in[D]$ a causal parent of $Y$ if $d\in S^{e,*}$ for at least one $e\in[E]$. Formally we define the set of causal parents of $Y$ as $S^{*}\vcentcolon=\bigcup_{e\in[E]}S^{e,*}$. Our inference goal of finding the causal parents of $Y$ is then equivalent to finding $S^{*}$.
3.1 Notation
This section introduces notation we will use throughout the paper. A graph $G=([D],\mathcal{E})$ is a tuple where indices in $[D]$ represent nodes and $\mathcal{E}\subseteq[D]^{2}$ represent directed edges between nodes, with the assumption that $(d,d)\not\in\mathcal{E}$ for any $d\in[D]$. If $(d_{1},d_{2})\in\mathcal{E}$ we call $d_{1}$ a parent of $d_{2}$ and we call $d_{2}$ a child of $d_{1}$. We call $d_{1}$ an ancestor of $d_{k}$ and we call $d_{k}$ a descendant of $d_{1}$ if there exists a sequence $d_{1},\cdots,d_{k}$ such that $(d_{i},d_{i+1})\in\mathcal{E}$ for $1\leq i<k$. If $d_{k}$ is not a descendant of $d_{1}$ we call $d_{k}$ a non-descendant of $d_{1}$. A node without any child is called a sink node. Given a node $d\in[D]$ we respectively define $\textbf{PA}(d)$, $\textbf{AN}(d)$, $\textbf{DE}(d)$, $\textbf{NDE}(d)$ as the set of all parents, ancestors, descendants and non-descendants of $d$.
Some further notation we use: With $X^{e}_{d,i}$ we indicate the entry of $X^{e}$ in the $d$-th row and $i$-th column, which corresponds to the $d$-th covariate of observation $i$. Similarly $\beta^{e}_{d}$ indicates the $d$-th entry of $\beta^{e}$. For $S\subseteq[D]$ we write $X^{e}_{S}\in\mathbb{R}^{n^{e}\times|S|}$ to indicate the sub-matrix of $X^{e}$ with columns given by $S$. In our notation the superscript $e$ always indicates the corresponding environment. We use $\mathbf{E}[\cdot]$, $\mathbf{V}[\cdot]$ and $\mathbf{C}[\cdot,\cdot]$ respectively as the expectation, variance and covariance operator. Finally we set $(\boldsymbol{X},\boldsymbol{Y})\vcentcolon=\{(X^{e},Y^{e})\}_{e\in[E]}$.
3.2 Data generation mechanism
Figure 1: The above graph shows the causal relations we assume, shown by the solid arrows. The dotted arrows may exist, but are not assumed. When we generate data from this type of model, we assume that we have access to $E$ different environments $e$ with structural parameter $\beta^{e}$ (outer box), and for each environment $e$ we observe $n^{e}$ samples according to Equation (1) (inner box).
While the previous subsections outline the main idea of our setting, we formalize it in the following assumption:
@@ 3.2 Data generation mechanism
There exists a subset $S^{*}\subseteq[D]$, a zero mean distribution $F^{*}$ and a collection of coefficients $\beta^{e}\in\mathbb{R}^{D}$ for $e\in[E]$, such that $\beta^{e}_{d}=0$ for all $d\not\in S^{*}$ and $e\in[E]$. Furthermore, for all $e\in[E]$ we have that
@@
At first, our setting may seem like a sparse regression task, but we emphasize that this model is considered to be a structural causal model, and the parameters $\beta^{e}$ have the interpretation of structural, rather than regression coefficients. We refer to Section 3.8.1 in Pearl (2016) for further discussion on this distinction. A particular consequence is that in condition (3) we require the unobserved noise $\varepsilon^{e}$ to be only independent of the causal parents $X^{e}_{S^{*}}$, but not necessarily of other variables. This becomes important when we want to find causal relations, instead of only statistical associations. Consider, for example, the simple causal structure of Figure 2. Here $X^{e}_{1}$ is a causal parent of $Y^{e}$ and independent of $\varepsilon^{e}$. However, $X^{e}_{2}$ is a descendant of $Y^{e}$ and thus not necessarily independent of $\varepsilon^{e}$.
We also draw special attention to condition (2), i.e., the assumption that $\varepsilon^{e}_{i}\sim F^{*}$ for all $e\in[E]$ and $i\in[n^{e}]$. This is arguably one of the stronger assumptions in our setting, and could be relaxed as discussed in Section 7. The main purpose of this assumption is, however, to ensure the environment $e$ does not have any influence on the unobserved variables. Considering the environment again as an intervention it becomes clear that some form of this assumption is necessary, as otherwise we have an unobserved confounding effect of the intervention.
It is also important to highlight that we do not make any independence assumptions across the different environments. The assumption that the distribution $F^{*}$ has zero mean is not crucial. This can always be achieved by adding an extra column to $X^{e}$, which is deterministically one, while the corresponding structural parameter in $\beta^{e}$ is then the negative mean of $F^{*}$.
Figure 2: In this simple model the variable $X^{e}_{2}$ is not necessarily independent of the noise $\varepsilon^{e}$.
Finally, we mention at several stages that the different environments should be heterogeneous, i.e., they should be different from each other in some aspect. In view of Assumption 1, heterogeneity corresponds either to different distributions of the covariates $X^{e}$, to different parameters $\beta^{e}$, or both. Within the context of our paper heterogeneity of environments is needed to identify causal parents. In Section 4.2, however, we provide sufficient conditions that ensure that the causal parents are identifiable for almost all possible values of $(\beta^{e})_{e\in[E]}$.
3.2.1 Examples
We now describe two settings in which Assumption 1 is sensible, so that we could apply LoLICaP, which we describe later in the paper.
Finding causal parents of a disease.
Assume we want to find causes for a certain disease. We collect data from different countries, which serve as the environments. We then collect data from plausible risk factors for the disease, where we have to make sure that we include all factors, which are heterogeneous across environments, such that the unexplained variation of the disease is roughly of the same distribution in all environments. Heterogeneous causal parents can be for example the diet, stress factors such as working conditions, lifestyle, genetic variations etc. In this setting LoLICaP is, in principle, able to identify the true causes of the disease amongst all plausible risk factors.
Finding causes with temporal data and mechanism shift.
@@ 3.2.1 Examples
We are collecting dynamical data from a machine and observe that a target variable $Y$ starts to drift, and we want to find the cause for that. (Note the subtle difference that we are now looking for the causal parents of the drift of $Y$, and not all causal parents of $Y$). The underlying, unknown, cause is that a certain component of the machine is degrading over time. Within our setting this corresponds, for example, to the parameter $\beta^{e}_{1}$ changing over time. We chose the environments as data from different time intervals, such that the local linearity condition is approximately true. In that case Assumption 1 is (approximately) fulfilled if time affects the drift only through the degradation of $\beta^{e}_{1}$.
3.3 Model strengths and limitations
As our prescribed default goal is to find all causal parents, a problem one may face is that some causal parents and the corresponding structural parameters are not heterogeneous enough across environments. Consider again the disease example and assume the diet was indeed a cause for the disease. If we collect data only from countries with the same diet and its structural parameter (i.e. the way the diet affects the disease) is always the same, LoLICaP will not find diet as a causal parent. From the perspective of environments as interventions it is clear that we are not intervening on the diet, and therefore can generally not draw causal conclusions about this covariate.
Another potential problem is that there is a causal parent which is heterogeneous in the environments, but which we do not observe. This corresponds to the violation of the assumption that the unobserved noise has a fixed distribution, i.e., $\varepsilon^{e}_{i}\sim F^{*}$ for all $e\in[E]$ and all $i\in[n^{e}].$ From the causal perspective this can be seen as the problem that the intervention is a confounder of the target.
@@ 4 On the identifiability of causal parents
Recall that, ultimately, we want to build a method based on observational data that is able to identify a set $\tilde{S}\subseteq[D]$ of causal parents. This set desirably approximates $S^{*}$ as well as possible, and ideally $\tilde{S}=S^{*}$. This section takes steps leading to a (meta)-methodology that ensures that $\tilde{S}\subseteq S^{*}$ with high probability, so that we guarantee at least control over false discoveries. This approach is rooted on a testing perspective, and aims to identify sets of plausible causal parents, which are also often called invariant sets (of covariates) in the relevant literature. Roughly speaking, a subset $S\subseteq[D]$ is plausible if it allows for a data generation model as described by Assumption 1, when $S$ takes the place of $S^{*}$. In that sense the name invariant set refers to the property that the covariates with indices in $S$ can model the target $Y$ with an additive noise $\varepsilon^{e}$ that is invariant in $e$. Importantly, the true parents $S^{*}$ are an invariant set by Assumption 1, so finding a minimal invariant set of some sorts allows us to control the false positive probability.

### v2

# Invariant Causal Prediction with Local Models

@@ Abstract
We consider the task of identifying the causal parents of a target variable among a set of candidates from observational data. Our main assumption is that the candidate variables are observed in different environments which may, under certain assumptions, be regarded as interventions on the observed system. We assume a linear relationship between target and candidates, which can be different in each environment with the only restriction that the causal structure is invariant across environments. Within our proposed setting we provide sufficient conditions for identifiability of the causal parents and introduce a practical method called L-ICP (Localized Invariant Causal Prediction), which is based on a hypothesis test for parent identification using a ratio of minimum and maximum statistics. We then show in a simplified setting that the statistical power of L-ICP converges exponentially fast in the sample size, and finally we analyze the behavior of L-ICP experimentally in more general settings.
@@ 1 Introduction
We consider the problem of identifying the causal parents of a target variable among a set of candidate variables, based only on observational data. As usual, causal inference and learning from observational data necessarily relies on assumptions. The main assumption used in this work is that data is collected in different environmental scenarios. An emblematic example is that of machine diagnostics, where we are monitoring several connected components of the machine. Different environments correspond, for example, to machines of the same model, but operating in different locations, different settings of a machine, or data collected in different points in time. If the system behaves differently across environments we talk about heterogeneous environments, and these can then be interpreted as accidental interventions on the system. The invariant causal prediction (ICP) principle, pioneered by Peters et al. (2016), is a particular way of using heterogeneous environments for causal discovery. It is based on the idea that the performance of any prediction model for the target variable should be invariant under interventions on the covariates, if and only if all covariates within the model are causal parents of the target. We extend upon that work by relaxing Peters et al. (2016) global linearity with a local linearity assumption, meaning that each environment is equipped with its own linear model. Relaxing the global model to local models carries a couple of interesting implications that we address in this paper. In particular, local models extend the scope of what might be considered an (accidental) intervention. While heterogeneity, as viewed in Peters et al. (2016), is always introduced by interventions on the covariate distributions, heterogeneous local models can be seen as informative interventions on the mechanisms within the system.
Consider, for example, the task of identifying key factors that influence fluctuations in stock prices and market volatility; this highlights the relevance of our extension. Within our framework, we first identify key times when important legislation concerning the stock market was enacted. We then choose our environments as the time intervals between these legislative actions. One may assume that a legislation intervenes on the distributions of the important factors, but also on the mechanism between those factors and stock prices/ market volatility. In our setting both types of (accidental) interventions are allowed and useful for causal discovery.
The paper is organized as follows: Section 2 contextualizes our contributions, positioning them relative to the related work. Section 3 formally introduces the setting, assumptions and our inference goals. In addition, it features further concrete examples showcasing the meaningfulness of the modeling assumptions. Section 4 introduces a meta-approach and characterizes sufficient conditions under which this approach can identify the causal parents of a target variable. Driven by that knowledge Section 5 introduces L-ICP, a practical approach to detect causal parents based on observational data. This approach inherits many of the properties of the idealized meta-procedure, and we show in a simplified setting that the power of L-ICP converges exponentially fast, in the sample size, towards one. Section 6 numerically illustrates the performance of L-ICP, and we conclude the paper in Section 7 with a discussion and an outlook on future work.
@@ 2 Related work
The idea of relying on heterogeneous environments for causal discovery is not new, and was instrumental for the work in Peters et al. (2016). This setup was also extended to sequential data in Pfister et al. (2019), while Heinze-Deml et al. (2017) investigate non-linear models by using conditional independence tests. Relaxations of the global linearity assumption towards local models have been also considered: Christiansen and Peters (2020) and Zhou et al. (2022) assume that the local structural parameters, i.e. the parameters of the assumed linear relationship between target and its parents, are related through a common (unobserved) variable. Huang et al. (2019, 2020) consider a temporal setting, where the local structural parameters change according to an auto-regressive model. In contrast to these works we consider a setting where the structural parameters can be radically different for the various environments.
The following works aggregate the information from different environments, without relying on heterogeneity assumptions, while rather non-Gaussian noise or non-linear relationship assumptions play a pivotal role: Osman et al. (2023) propose a method that finds causal structures and detects interventions using a minimum description length score on the causal factorization, which they define jointly over the different environments. They assume non-linear relationships, identifiability of the DAG up the Markov equivalence class and a low noise condition. Chen et al. (2021) assume a linear structural equation model (SEM), such that at least in some environments the corresponding DAG is identifiable from data. Shimizu (2012) proposes an extension of LiNGAM (Shimizu et al., 2006) to the multiple environment case. LiNGAM is a method that can find DAGs under a linearity and non-Gaussianity assumption. Finally, Mooij et al. (2020) developed the joint causal inference (JCI) framework, where environments (in the paper called contexts) are directly encoded as part of the structural causal model (SCM).
@@ 3 Setting
In this work we consider a scenario with two observable quantities of interest: a target, denoted by $Y$, and a set of covariates $X\vcentcolon=(X_{1},\cdots,X_{D})$ for $D\in\mathbb{N}$. Our overarching goal is to identify which of the $D$ covariates are the causal parents of the target $Y$. We further assume to have access to $E\in\mathbb{N}$ different environments and in each environment we receive $n^{e}$ observations, so that $Y^{e}\in\mathbb{R}^{n^{e}}$ and $X^{e}\in\mathbb{R}^{n^{e}\times D}$ are respectively the target and covariate observations. With $X^{e}_{d,i}$ we indicate the entry of $X^{e}$ in the $d$-th row and $i$-th column, which corresponds to the $d$-th covariate of observation $i$. For $S\subseteq[D]$ we write $X^{e}_{S}\in\mathbb{R}^{n^{e}\times|S|}$ to indicate the sub-matrix of $X^{e}$ with columns given by $S$. We assume that for each $e\in[E]$ the structural equation of $Y^{e}$ is given by
$Y^{e}\vcentcolon=X^{e}\beta^{e}+\varepsilon^{e},$ (1)
where $\varepsilon^{e}\in\mathbb{R}^{n^{e}}$ is a zero-mean random perturbation (specified explicitly below) and $\beta^{e}\in\mathbb{R}^{D}$ is the column vector of structural parameters. In the following $\beta^{e}_{d}$ indicates the $d$-th entry of $\beta^{e}$. As we consider the relationship $Y^{e}\vcentcolon=X^{e}\beta^{e}+\varepsilon^{e}$ to be a structural causal model (SCM) in the sense of Pearl (2016), we consider the set $S^{e,*}\vcentcolon=\{d\in[D]\mid\beta^{e}_{d}\neq 0\}\subseteq[D]$ as the true causal parents of $Y^{e}$. Correspondingly we define the set of causal parents of $Y$ as $S^{*}\vcentcolon=\bigcup_{e\in[E]}S^{e,*}$. Our inference goal of finding the causal parents of $Y$ is then equivalent to finding $S^{*}$. It is important to note the vectors $\beta^{e}$ can be radically different across environments. The following assumption formalizes our setting and further specifies the independence assumptions made:
@@
Let $S^{*}\subseteq[D]$ be defined as above, such that $\beta^{e}_{d}=0$ for all $d\not\in S^{*}$ and $e\in[E]$. There exists a zero mean distribution $F^{*}$ such that for all $e\in[E]$ we have that
@@
Finding causal parents of a disease. Assume we want to find causes for a certain disease with data from different countries, playing the role of different environments. We collect data from plausible risk factors for the disease (e.g., diet, lifestyle, genetic variations, etc.). Very plausibly risk factors are heterogeneous across different countries, and thus provides a scenario in which our setting and the resulting methodology are applicable. Next to the heterogeneity in the risk factors, unobserved factors such as the quality of the health care may also introduce heterogeneity in the structural parameters: a negative health outcome, given all risk factors, is much more likely if the health care is poor. In that case it becomes necessary to have local models.
@@
Finding causes of a mechanism shift. We are collecting dynamical data from a machine and observe that a target variable $Y$ starts to drift, and we want to find the cause for that. The underlying, unknown, cause is that a certain component of the machine is degrading over time. This degradation naturally provides heterogeneous environments, if we set our environments as different time intervals of the observations. Note that in this setting (2) is deemed quite reasonable and might embody sensor measurement noise.
3.1 Additional Notation
Some further notation we use throughout the paper: A graph $G=([D],\mathcal{E})$ is a tuple where indices in $[D]$ represent nodes and $\mathcal{E}\subseteq[D]^{2}$ represent directed edges between nodes, with the assumption that $(d,d)\not\in\mathcal{E}$ for any $d\in[D]$. If $(d_{1},d_{2})\in\mathcal{E}$ we call $d_{1}$ a parent of $d_{2}$ and we call $d_{2}$ a child of $d_{1}$. We call $d_{1}$ an ancestor of $d_{k}$ and we call $d_{k}$ a descendant of $d_{1}$ if there exists a sequence $d_{1},\cdots,d_{k}$ such that $(d_{i},d_{i+1})\in\mathcal{E}$ for $1\leq i<k$. If $d_{k}$ is not a descendant of $d_{1}$ we call $d_{k}$ a non-descendant of $d_{1}$. A node without any child is called a sink node. Given a node $d\in[D]$ we respectively define $\textbf{PA}(d)$, $\textbf{AN}(d)$, $\textbf{DE}(d)$, $\textbf{NDE}(d)$ as the set of all parents, ancestors, descendants and non-descendants of $d$. Furthermore, we use $\mathbf{E}[\cdot]$, $\mathbf{V}[\cdot]$ and $\mathbf{C}[\cdot,\cdot]$ respectively as the expectation, variance and covariance operator. Finally we set $(\boldsymbol{X},\boldsymbol{Y})\vcentcolon=\{(X^{e},Y^{e})\}_{e\in[E]}$.
@@ 4 On the identifiability of causal parents
Ultimately, our goal is to identify, based on data, a set $\tilde{S}\subseteq[D]$ of variables deemed causal parents, that is ideally identical to $S^{*}$. Towards this goal we develop a test-based methodology ensuring $\tilde{S}\subseteq S^{*}$ with high probability. The methodology works by identifying sets of plausible causal parents, which are also often called invariant sets (of covariates) in the relevant literature. Roughly speaking, a subset $S\subseteq[D]$ is plausible if it allows for a data generation model as described by Assumption 1, when $S$ takes the place of $S^{*}$. The inferred set of causal parents $\tilde{S}$ is then the intersection of all plausible sets. In the following section we formalize those concepts and show that we can control the false positive discoveries, so that $\tilde{S}\subseteq S^{*}$ with high probability.